# Simple and Provable Scaling Laws for the Test-Time Compute of Large Language Models

### v2

@@ 4.2 Results for the league-style algorithm
Figure 6 shows that the accuracy of the league-style algorithm grows with $N$ until saturation, for both datasets and all of Llama3.1, Qwen2.5 and Mixed options. For example, accuracy of Mixed shows a $8\%$ improvement on GPQA (from $45\%$ to $53\%$) as $N$ grows from 1 to 16, and a $4\%$ improvement (from $72\%$ to $76\%$) on MMLU-Pro-S. The Mixed option consistently outperforms Llama3.1 and Qwen2.5, similar to the case for the knockout-style algorithm. Given the same initial candidates, the league-style algorithm achieves higher accuracy than the knockout-style algorithm (dotted lines) does in some cases and lower in other cases, although the differences are minor in general.
Figure 6: Accuracy versus the number of initial candidates $N$ for the league-style algorithm (LG, solid lines), as well as for the knockout-style algorithm (KO, dotted lines) given the same initial candidates.
Connection to Assumption 2.
Figure 7: The distribution of problems from the GPQA (top) or MMLU-Pro-S (bottom) benchmark, characterized by $\widehat{p}_{\text{gen}}$ and $\widehat{\Delta}$ that are estimated with the empirical results for the league-style algorithm using the Llama3.1 (left), Qwen2.5 (middle) or Mixed (right) option. Each plot is annotated with the number of problems satisfying the condition $\widehat{\Delta}>0$, $\widehat{\Delta}\leq 0$, $\widehat{p}_{\text{gen}}=0$ or $\widehat{p}_{\text{gen}}=1$. To the right of each plot is a histogram for $\widehat{\Delta}$. Each problem is represented by a circle if it is solved correctly by the league-style algorithm with $N=16$, and by a cross otherwise. We neglect problems with $\widehat{p}_{\text{gen}}=0$ or $1$, i.e., problems for which the initial candidate solutions are all incorrect or all correct, since there is no way of obtaining meaningful estimate of $\widehat{\Delta}$ for such problems.
Figure 8: Accuracy versus the number of initial candidates $N$ for the league-style algorithm, on a filtered subset of problems satisfying $10<\widehat{p}_{\text{gen}}<$ 1 and $\widehat{\Delta}>0$. The values of accuracy are calculated with new trials of the algorithm, which are statistically independent of the trials used for estimating $\widehat{p}_{\text{gen}}$ and $\widehat{\Delta}$.
@@
We conduct further analysis that draws connections to the theories developed in Section 3 and enhances our understanding of the empirical performance of the league-style algorithm. Let us start by finding proxies for the $p_{\text{gen}}$ and $\Delta$ parameters in Assumption 2 for each problem of both datasets. For the former, we define $\widehat{p}_{\text{gen}}$ again as the fraction of initial candidate solutions with the correct final answer. For the latter, we define $\widehat{\Delta}\coloneqq\widehat{\mu}_{i_{1}}-\widehat{\mu}_{i_{2}}$, where $i_{1}\coloneqq\arg\max_{i\in[N]:\ y_{i}\text{ is correct}}\widehat{\mu}_{i}$ and $i_{2}\coloneqq\arg\max_{i\in[N]:\ y_{i}\text{ is incorrect}}\widehat{\mu}_{i}$ are the indices for the strongest correct candidate and for the strongest incorrect candidate, respectively. Note that by definition, for any problem with $\widehat{p}_{\text{gen}}\notin\{0,1\}$, the league-style algorithm returns a correct solution to the problem if and only if $\widehat{\Delta}>0$.
@@
Figure 7 characterizes the distribution of problems from the GPQA and MMLU-Pro-S datasets, in terms of $\widehat{p}_{\text{gen}}$ and $\widehat{\Delta}$. It is noteworthy that there exists a non-trivial proportion of problems for which $\widehat{p}_{\text{gen}}$ is fairly small (i.e., most of the initial candidates are incorrect), while the league-style aggregation stage manages to attain a positive $\widehat{\Delta}$ and thus identify a correct candidate for the final output, which demonstrates the benefits of the league-style aggregation method. We also pay special attention to the subset of problems satisfying $10<\widehat{p}_{\text{gen}}<$ 1 and $\widehat{\Delta}>0$, which correspond to the conditions stated in Assumption 2, except that those easy problems with $\widehat{p}_{\text{gen}}=1$ are excluded. We run a new, independent trial of the league-style algorithm (with Llama3.1/Qwen2.5/Mixed) on this subset. Figure 8 confirms that accuracy improves with $N$ significantly on this subset (e.g., from 57% to 82% for Mixed), as predicted by our theories.
@@
While all previous experiments consider the round-robin version of the league-style algorithm, we also wonder if it is feasible to improve its computational efficiency by estimating the average win rate of each candidate with $M<N-1$ subsampled opponents, while maintaining its accuracy. The empirical results in Figure 9 provide a positive answer and match what our theories in Section 3 predict: (i) accuracy initially increases with $M$, which confirms the benefits of comparing each candidate with multiple opponents; (ii) once $M$ exceeds a threshold (around 4 or 5) that is much smaller than $N=16$, accuracy saturates around the level achieved by the round-robin version, but at a lower computational cost.
Figure 9: Accuracy versus $M$, the number of subsampled opponents for each candidate, for the league-style algorithm with $N=16$.
@@ 5 Related works
There has been a surge of interest in enhancing problem solving with LLMs by various strategies of scaling up test-time compute, including those mentioned earlier in Section 1. Extensive literature review on this topic can be found in some recent surveys [39, 47, 11], and thus we will focus our discussions in this section mainly on prior works that are closely related to ours.
A provable inference scaling law (in the sense of Definition 1) has been developed for majority voting in the literature [3, 42], which essentially requires two assumptions. (i) First, it should be feasible to divide the candidate solutions into several groups and have a meaningful count for each group. This is true for, say, multiple-choice questions, but not for many other tasks where all candidate solutions are distinct, such as open-ended writing. In contrast, both knockout-style and league-style algorithms proposed in our work, where selecting a good solution out of many candidates is reduced to a series of pairwise comparisons (or two-choice questions), can be applied to a wider range of tasks in a unified manner. (ii) Second, majority voting requires that the probability of generating a solution that belongs to the correct group, akin to $p_{\text{gen}}$ in our notations, should be larger than that of any other group; otherwise, it has been proved that the success probability of majority voting will converge to zero instead as its test-time compute grows [3]. In comparison, our proposed algorithms only require $p_{\text{gen}}>0$, though with additional assumption about LLMs’ capabilities of conducting pairwise comparisons.
Deriving a provable scaling law for the Best-of-N approach is straightforward, provided that a perfect verifier is available for picking a correct solution out of $N$ candidates: if one LLM call generates a correct solution with probability at least $p_{\text{gen}}>0$, then the failure probability of Best-of-N is upper bounded by $(1-p_{\text{gen}})^{N}$, which converges to zero as the number of samples $N$ grows. One obvious limitation, of course, is that verifiers are unavailable or imperfect in many practical scenarios, which can hinder the performance of Best-of-N as its test-time compute grows [5, 32].
@@
Our algorithm design in this work has drawn inspiration from various areas. For example, the essential idea of pairwise comparison has been prominent in LLM alignment [1, 27, 29], as well as in the LLM-as-a-judge approach [51, 19] that has been used mostly for the purpose of scalable and automatic LLM evaluation, but also as a key component of test-time strategies in some prior works (cf. Section 5.4 of [19]). Although it is possible to verify, score or refine a solution by itself [13, 24, 7, 10], intuitively it is often much easier (for LLMs or human) to detect the errors or hallucinations in an incorrect solution when it is placed right next to a correct one, or evaluate the quality of a solution by comparing it to another one. The knockout and league tournaments have also been investigated in prior LLM research [16, 22, 50, 18, 12], albeit with purposes or concrete implementations that are different from ours. Given this context, we remark that the main novelty and contributions in this work are perhaps less about the proposed two-stage algorithms themselves, but rather more about developing rigorous and theoretical understanding of their underlying assumptions and efficacy (via clearly identifying sufficient conditions for boosting their success probability up to $100\%$ and formally deriving quantitative bounds for their computational and sample complexities), and demonstrating their promising empirical performance through extensive experiments.
With the ongoing advancements of LLMs (such as OpenAI’s o1 [26, 47]), we anticipate that stronger LLMs in the future will achieve higher values of $p_{\text{gen}}$, $p_{\text{comp}}$ and $\Delta$ in Assumptions 1 and 2 for the same input problem, which will reduce the number of LLM calls required by our algorithms, according to the proposed theorems. Furthermore, these technical assumptions will become valid for more challenging problems that currently violate them (due to limitations of the existing LLMs), which will enable the proposed algorithms to achieve success in broader scenarios.
6 Discussions
We conclude by discussing some limitations of the current work as well as directions for future work.
Solving hard problems by task decomposition.
Our theorems are not directly applicable to challenging scenarios where the LLM simply has zero chance of generating a correct solution to the input problem in one shot, i.e., when $p_{\text{gen}}=0$. For such scenarios, a generic approach adopted by practical LLM-based algorithms and agentic workflows [4, 46, 43] is to decompose the original problem into smaller sub-problems and solve each of them. As long as the sub-problems satisfy the technical assumptions in our theories, we can apply the proposed algorithms to each of them. This approach can offer not only formal guarantees for the success probability of solving the original problem, but also higher efficiency compared to solving it directly.
@@ 6 Discussions
To see this, consider a scenario where solving the original problem requires solving all $S\geq 1$ sub-problems correctly, and each sub-problem satisfies Assumption 1 with parameters $p_{\text{gen}}>0$ and $p_{\text{comp}}>0.5$. Directly solving all $S$ sub-problems has a exponentially small success probability $p_{\text{gen}}^{S}$, and thus generating a correct solution alone already requires $\Omega((1/p_{\text{gen}})^{S})$ attempts, not to mention identifying which attempt is successful. In contrast, by conducting task decomposition and applying the knockout-style algorithm (with hyperparameters $N$ and $K$) to each sub-problem, an overall success probability of $1-\delta$ for solving the original problem can be guaranteed as long as the failure probability for each sub-problem is bounded by $\delta/S$, thanks to the union bound. According to Eq. (5), this is guaranteed with
@@
Note that $N$ and $K$ only have logarithmic dependence on the number of sub-problems $S$. The total number of LLM calls with this approach is $(K+1)\times N\times S$ (cf. Section 2.2), which grows with $S$ linearly, up to logarithmic factors.
More general problem settings.
One limitation of the current work is that, we have focused on the binary metric of correctness throughout, and conducted experiments with only two benchmarks (due to limited computational resources). Future work may extend the theories to more general performance metrics, and test the proposed algorithms in broader scenarios.
Another different problem setting that we have not investigated is about optimal allocation of test-time compute [31, 3, 6, 48]: given a fixed budget of test-time compute and a set of problems to be solved, is there a method for automatically finding the optimal configuration for each problem (e.g., the values of $N$ and $K$ in our algorithms), in order to maximize the overall performance, such as the number of solved problems? Achieving this for the proposed algorithms will likely be a fruitful direction for future research.
Finally, let us conclude with some discussions about anytime algorithms [40]. In many real-world scenarios, the available amount of test-time compute is adaptive and unknown a priori. To address such cases, we can easily convert the knockout-style algorithm to an “anytime” variant that does not require pre-specifying $N$. For example, the algorithm might start with 4 candidate solutions and choose the winner via a knockout tournament. If more test-time compute is allowed (e.g., the user is not eagerly requesting the solution, or more computational resources become available), then the algorithm can launch another tournament with 4 freshly sampled candidates, the winner of which will compete with the winner of the previous tournament. This complete process is indeed equivalent to a single tournament with $N=4+4=8$. Such a process can continue until the user finally requests the solution; the eventual value of $N$ is determined online and automatically achieves the maximum value allowed by the available test-time compute. Similarly, the league-style algorithm can be converted to an anytime variant, where the total number of candidates and/or the number of comparisons for each candidate increase gradually as more test-time compute becomes available. It would be interesting future work to investigate such anytime algorithms from a theoretical perspective, or apply these methods to solving practical problems in online scenarios.
@@ A.3 Proof of Theorem 3
${\mathbb{P}}\Big{(}y_{i}\notin\mathcal{Y}_{\text{cs}},\forall i\in[N]\Big{)}\leq(1-p_{\text{gen}})^{N}.$
@@
Note that in the last line, $y_{o_{i}(j)}\sim\text{Unif}(y_{j},j\in[N]\backslash\{i\})$, and $r_{i}(j)\sim\mathcal{M}_{\text{comp}}(x,y_{i},y_{o_{i}(j)})$. By Hoeffding’s inequality, we have the following for a specific $i\in[N]$:
@@
${\mathbb{P}}\Big{(}|\widetilde{\mu}_{i}-\mu_{i}|\geq\frac{\Delta}{4}\Big{)}$ $={\mathbb{P}}\bigg{(}\Big{|}\frac{1}{N-1}\sum_{j\in[N]\backslash\{i\}}{\mathbb{E}}_{r\sim\mathcal{M}_{\text{comp}}(x,y_{i},y_{j})}\phi(r,y_{i},y_{j})-\mu_{i}\Big{|}\geq\frac{\Delta}{4}\bigg{)}\leq 2\exp\Big{(}-\frac{(N-1)\Delta^{2}}{8}\Big{)}.$
@@
${\mathbb{P}}\Big{(}|\widehat{\mu}_{i}-\mu_{i}|\geq\frac{\Delta}{2}\Big{)}\leq{\mathbb{P}}\Big{(}|\widehat{\mu}_{i}-\widetilde{\mu}_{i}|\geq\frac{\Delta}{4}\Big{)}+{\mathbb{P}}\Big{(}|\widetilde{\mu}_{i}-\mu_{i}|\geq\frac{\Delta}{4}\Big{)}\leq 2\exp\Big{(}-\frac{K\Delta^{2}}{8}\Big{)}+2\exp\Big{(}-\frac{(N-1)\Delta^{2}}{8}\Big{)}.$
@@
$1-(1-p_{\text{gen}})^{N}-2N\exp\Big{(}-\frac{K\Delta^{2}}{8}\Big{)}-2N\exp\Big{(}-\frac{(N-1)\Delta^{2}}{8}\Big{)},$

### v3

@@ 4.2 Results for the league-style algorithm
Figure 5 shows that the accuracy of the league-style algorithm grows with $N$ for all of Llama3.1, Qwen2.5 and Mixed options on GPQA, e.g., accuracy of Mixed improves by 8 points (from $45\%$ to $53\%$) as $N$ grows to 16. The Mixed option consistently outperforms Llama3.1 and Qwen2.5, similar to the case for the knockout-style algorithm. Given the same initial candidates, the league-style algorithm achieves higher accuracy than the knockout-style algorithm does in some cases and lower in other cases, although the differences are minor in general.
Bridging with the theories.
@@
Let us start by finding proxies for the $p_{\text{cs}}$ and $\Delta$ parameters in Assumption 2. For the former, we define $\widehat{p}_{\text{cs}}$ as the fraction of initial candidate solutions with the correct final answer. For the latter, we define $\widehat{\Delta}\coloneqq\widehat{\mu}_{i_{1}}-\widehat{\mu}_{i_{2}}$, where $i_{1}\coloneqq\arg\max_{i\in[N]:\ y_{i}\text{ is correct}}\widehat{\mu}_{i}$ and $i_{2}\coloneqq\arg\max_{i\in[N]:\ y_{i}\text{ is incorrect}}\widehat{\mu}_{i}$ are the indices for the strongest correct candidate and for the strongest incorrect candidate, respectively. Note that by definition, for any problem with $\widehat{p}_{\text{cs}}\notin\{0,1\}$, the league-style algorithm returns a correct solution to the problem if and only if $\widehat{\Delta}>0$.
@@
Figure 15 in Appendix C.3 characterizes the distribution of GPQA and MMLU-Pro-S problems in terms of $\widehat{p}_{\text{cs}}$ (X-axis) and $\widehat{\Delta}$ (Y-axis); the plot corresponding to GPQA and Mixed can also be found in Figure 5. It is noteworthy that there exists a non-trivial proportion of problems for which $\widehat{p}_{\text{cs}}$ is fairly small (i.e., most of the initial candidates are incorrect), yet the proposed league-style aggregation stage still manages to attain a positive $\widehat{\Delta}$ and thus identify a correct candidate for the final output. Figure 5 also confirms that significant improvements in accuracy, e.g., a 25% increase for Mixed, can be achieved on the subset of problems that approximately satisfy Assumption 2.
@@
While all previous experiments consider the round-robin version of the league-style algorithm, we also wonder if it is feasible to improve its computational efficiency by estimating the average win rate of each candidate with $M<N-1$ subsampled opponents, while maintaining its accuracy. The empirical results in Figure 5 provide a positive answer and match what our theories in Section 3 predict: (1) accuracy initially increases with $M$, which confirms the benefits of comparing each candidate with multiple opponents; (2) once $M$ exceeds a threshold (around 4 or 5) that is much smaller than $N=16$, accuracy saturates around the level achieved by the round-robin version, but at a lower computational cost.
Figure 5: Empirical results for the league-style algorithm. (1) Accuracy versus the number of initial candidates $N$ for the league-style (LG, solid lines) and knockout-style (KO, dotted lines) algorithms, given the same initial candidates. (2) The distribution of GPQA problems, characterized by $\widehat{p}_{\text{cs}}$ and $\widehat{\Delta}$ that are estimated with the Mixed option. Each problem is represented by a circle if it is solved correctly at $N=16$, and by a cross otherwise. (3) Accuracy versus $N$ on the subset of problems satisfying $10<\widehat{p}_{\text{cs}}<$ 1 and $\widehat{\Delta}>0$. The values of accuracy are calculated with new trials of the algorithm, thus statistically independent of $\widehat{p}_{\text{cs}}$ and $\widehat{\Delta}$. (4) Accuracy versus $M$, the number of subsampled opponents for each candidate, for the league-style algorithm with $N=16$.
@@ 5 Related works
There exist other test-time strategies that enjoy provable inference scaling laws in the sense of Definition 1. One example is majority voting, whose provable success requires two assumptions [3, 48]: (1) it is feasible to divide the candidate solutions into several groups and have a meaningful count for each group (which is not the case in tasks like open-ended writing, where all candidate solutions are distinct), and (2) the probability that one LLM call generates a solution belonging to the correct group is higher than that for any other group. In comparison, our proposed algorithms are free from the first restriction, and only require $p_{\text{gen}}>0$ while making additional assumptions about LLMs’ capabilities in pairwise comparisons. Another example is Best-of-N (BoN) sampling, for which deriving a provable scaling law is straightforward provided that a perfect verifier is available: if one LLM call generates a correct solution with probability $p_{\text{gen}}>0$, then the failure probability of BoN is $(1-p_{\text{gen}})^{N}$. One obvious limitation is that verifiers are unavailable or imperfect in many practical scenarios, which can hinder the performance of BoN [5, 37, 2]. We refrain from comparing our methods with BoN in our experiments, since introducing an external verifier or reward model will bring extra variability that makes it difficult to conduct a fair and meaningful empirical comparison.
@@
Our algorithm design has drawn inspiration from various areas. For example, the essential idea of pairwise comparison has been prominent in LLM alignment [1, 30, 33] and the LLM-as-a-judge paradigm [57, 20]. Although it is possible to verify, score or refine a solution by itself [14, 25, 6, 11], it is often much easier (for LLMs or human) to detect the errors or hallucinations in an incorrect solution when it is placed right next to a correct one, or evaluate the quality of a solution by comparing it to another one. The knockout and league tournaments have also been investigated in prior LLM research [17, 23, 56, 19, 13], albeit with purposes or implementations that are different from ours. Given this context, we remark that the main novelty and contributions in this work are perhaps less about the proposed two-stage algorithms themselves, but rather more about developing rigorous and theoretical understanding of their underlying assumptions and efficacy (via clearly identifying sufficient conditions for boosting their success probability up to $100\%$ and formally deriving quantitative bounds for their computational and sample complexities), and demonstrating their promising empirical performance through extensive experiments.
6 Limitations and future work
One limitation of this work is that, like any other test-time scaling method, the proposed algorithms trade computation for a higher success rate. Future work may try to find practical ways to determine the smallest values of hyperparameters $N$ and $K$ necessary for a targeted success probability. We also note that the provable success of the proposed algorithms relies on technical assumptions that might not hold true in the first place (as is the case for most theories), although we anticipate optimistically that with the ongoing developments of LLMs, the assumptions made in this work (and thus our algorithms) will automatically become feasible for more and more challenging tasks. Future work may also try to extend the methodologies and theories to broader scenarios, some of which are elaborated in the following.
Combination with other test-time strategies.
The proposed two-stage methods are orthogonal and complementary to many other test-time scaling strategies. For example, our experiment with QwQ-32B shows that the performance of a long-CoT reasoning LLM can be further boosted with the knockout-style algorithm. Future work may investigate systematic approaches of combining different inference scaling strategies for achieving the best performance [36].
Application to agentic workflows.
@@ 6 Limitations and future work
In complex real-world scenarios, a common practice is to adopt an agentic workflow that decomposes the original task into manageable sub-problems and involves multiple LLM calls to solve all of them [52, 49, 4]. Applying the knockout/league-style algorithm proposed in this work to each sub-problem can efficiently amplify the success probability of the overall workflow. To see how this works, consider a scenario where solving the original problem requires solving all $S\geq 1$ sub-problems correctly, and each sub-problem satisfies Assumption 1 with parameters $p_{\text{gen}}>0$ and $p_{\text{comp}}>0.5$. Directly solving all $S$ sub-problems has a exponentially small success probability $p_{\text{gen}}^{S}$, and thus generating a correct solution alone already requires $\Omega((1/p_{\text{gen}})^{S})$ attempts, not to mention identifying which attempt is successful. In contrast, by applying the knockout-style algorithm (with hyperparameters $N$ and $K$) to each sub-problem, an overall success probability $1-\delta$ for solving the original problem can be guaranteed as long as the failure probability for each sub-problem is bounded by $\delta/S$, thanks to the union bound. According to Eq. (1), this is guaranteed with
@@
both with logarithmic dependence on $S$. The total number of LLM calls with this approach is $(K+1)\times N\times S$ (cf. Section 2.2), which grows with $S$ linearly, up to logarithmic factors.
Anytime algorithms for online settings.
In many real-world scenarios, the available amount of test-time compute is adaptive and unknown a priori. To address such cases, we can easily convert the knockout-style algorithm to an “anytime” variant [46] that does not require pre-specifying $N$. For example, the algorithm might start with 4 candidate solutions and choose the winner via a knockout tournament. If more test-time compute is allowed (e.g., the user is not eagerly requesting the solution, or more computational resources become available), then the algorithm can launch another tournament with 4 freshly sampled candidates, the winner of which will compete with the winner of the previous tournament. This complete process is indeed equivalent to a single tournament with $N=4+4=8$. Such a process can continue until the user finally requests the solution; the eventual value of $N$ is determined online and automatically achieves the maximum value allowed by the available test-time compute. Similarly, the league-style algorithm can be converted to an anytime variant, where the total number of candidates and/or the number of comparisons for each candidate increase gradually as more test-time compute becomes available. It would be interesting future work to investigate such anytime algorithms from a theoretical or practical perspective.
@@ A.3 Proof of Theorem 3
${\mathbb{P}}\Big{(}y_{i}\notin\mathcal{Y}_{\text{cs}},\forall i\in[N]\Big{)}\leq(1-p_{\text{cs}})^{N}.$
@@
Note that in the last line, $y_{o_{i}(j)}\sim\text{Unif}(y_{j},j\in[N]\backslash\{i\})$, and $r_{i}(j)\sim\mathcal{M}_{\text{comp}}(x,y_{i},y_{o_{i}(j)})$. By Hoeffding’s inequality, we have the following for each $i\in[N]$:
@@
${\mathbb{P}}\Big{(}|\widetilde{\mu}_{i}-\mu_{i}|\geq\frac{\Delta}{4}\Big{)}$ $={\mathbb{P}}\bigg{(}\Big{|}\frac{1}{N-1}\sum_{j\in[N]\backslash\{i\}}{\mathbb{E}}_{r\sim\mathcal{M}_{\text{comp}}(x,y_{i},y_{j})}\phi(r,y_{i},y_{j})-\mu_{i}\Big{|}\geq\frac{\Delta}{4}\bigg{)}$
$\leq 2\exp\Big{(}-\frac{(N-1)\Delta^{2}}{8}\Big{)}.$
@@
${\mathbb{P}}\Big{(}|\widehat{\mu}_{i}-\mu_{i}|\geq\frac{\Delta}{2}\Big{)}$ $\leq{\mathbb{P}}\Big{(}|\widehat{\mu}_{i}-\widetilde{\mu}_{i}|\geq\frac{\Delta}{4}\Big{)}+{\mathbb{P}}\Big{(}|\widetilde{\mu}_{i}-\mu_{i}|\geq\frac{\Delta}{4}\Big{)}$
$\leq 2\exp\Big{(}-\frac{K\Delta^{2}}{8}\Big{)}+2\exp\Big{(}-\frac{(N-1)\Delta^{2}}{8}\Big{)}.$
@@
$1-(1-p_{\text{cs}})^{N}-2N\exp\Big{(}-\frac{K\Delta^{2}}{8}\Big{)}-2N\exp\Big{(}-\frac{(N-1)\Delta^{2}}{8}\Big{)},$